# Human2Robot: Learning Robot Actions from Paired Human-Robot Videos

### v3

@@ Abstract
Distilling knowledge from human demonstrations is a promising way for robots to learn and act. Existing methods, which often rely on coarsely-aligned video pairs, are typically constrained to learning global or task-level features. As a result, they tend to neglect the fine-grained frame-level dynamics required for complex manipulation and generalization to novel tasks. We posit that this limitation stems from a vicious circle of inadequate datasets and the methods they inspire. To break this cycle, we propose a paradigm shift that treats fine-grained human-robot alignment as a conditional video generation problem. To this end, we first introduce H&R, a novel third-person dataset containing 2,600 episodes of precisely synchronized human and robot motions, collected using a VR teleoperation system. We then present Human2Robot, a framework designed to leverage this data. Human2Robot employs a Video Prediction Model to learn a rich and implicit representation of robot dynamics by generating robot videos from human input, which in turn guides a decoupled action decoder. Our real-world experiments demonstrate that this approach not only achieves high performance on seen tasks but also exhibits significant one-shot generalization to novel positions, objects, instances, and even new task categories.
@@ 1 Introduction
The ability to observe others, whether humans or animals, and acquire skills to solve new tasks is a fundamental reason why humans excel at tackling a wide range of complex challenges. To enable robots to assist with diverse real-world problems, it is essential that they develop a similar capability, particularly the ability to learn directly from human demonstrations. This has motivated extensive research (Bahl et al. 2023; Srirama et al. 2024; Nair et al. 2022; Wang et al. 2023; Bahl, Gupta, and Pathak 2022; Smith et al. 2019) on how to learn from human-provided demonstrations effectively.
Despite recent progress, current approaches face a fundamental generalization gap. While they perform well on tasks encountered during training, they often fail entirely when presented with human demonstrations of unseen tasks. The prevailing paradigm relies on applying self-supervised (Xu et al. 2023b) or contrastive learning (Jain et al. 2024) methods to coarsely aligned human-robot video pairs. However, the lack of fine-grained supervision in existing datasets fundamentally constrains what models can learn. Rather than uncovering detailed action-level correspondences, models are often limited to capturing global features or high-level task summaries. Consequently, many approaches extract holistic video representations using models such as the Perceiver Resampler (Alayrac et al. 2022), which compresses entire clips into fixed-length embeddings. This severely limits their ability to model the nuanced frame-level temporal dynamics that are essential for generalization.
@@
In this paper, we propose that the key to unlocking fine-grained human-robot alignment lies in video generation (Ho et al. 2022; Esser et al. 2023; Blattmann et al. 2023a; Wu et al. 2023; Khachatryan et al. 2023; Xing et al. 2024; Tu et al. 2024a). Instead of merely learning whether two videos are broadly similar or mapping task goals from the human domain to the robot domain, we aim to generate a corresponding robot video directly from a human demonstration. We posit that by training a model to predict the precise frame-by-frame evolution of a robot’s movements, it learns a far richer and more temporally coherent alignment. This generative process forces the model to internalize the intricate dynamics of manipulation, enabling it to understand how a task is done, not just what is done.
@@ 2 Related Work
Teleoperation. Recently, VR-based methodologies (Iyer et al. 2024; Ding et al. 2024) have attracted considerable attention due to their cost-effectiveness, efficiency and versatility. However, these methods focus on controlling the robot, whereas our goal of using VR is to capture perfectly aligned videos of humans and robots, which are essential for robotic imitation.
Learning from Human Videos. Researchers have recently been attempting to leverage existing human-centric video datasets to enhance robot policy learning (Liu et al. 2018; Smith et al. 2020; Chen, Nair, and Finn 2021; Shaw, Bahl, and Pathak 2022; Zeng et al. 2024). Researchers propose to learn representations from human videos to assist in task execution (Xiao et al. 2022; Wang et al. 2023; Majumdar et al. 2023). However, these approaches need substantial prior knowledge and struggle to transfer to robots (Nair et al. 2022; Bahl et al. 2023; Bahl, Gupta, and Pathak 2022). Meanwhile, some human-video-conditioned methods  (Bharadhwaj et al. 2024; Jain et al. 2024; Xu et al. 2023b) have focused on aligning representations across human and robot videos, they are neither efficient nor capable of generalization. In contrast, our approach uses paired data and diffusion models to achieve strong generalization capabilities.
Diffusion Models for Video Generation. Current research has achieved remarkable performance of video generation (Ho et al. 2022; Esser et al. 2023; Blattmann et al. 2023a; Wu et al. 2023; Khachatryan et al. 2023; Xing et al. 2024; Tu et al. 2024a, b, c). Moreover, diffusion models for video stylization (Ye et al. 2025; Liu et al. 2023) and human image animation (Hu 2024; Tu et al. 2024a) have also shown its promising potential for visual and motion transfer. Moreover, diffusion model not only performs impressively in generative domains but also shows promising applications in visual understanding task. Some notable approaches involve using diffusion models for object detection (Chen et al. 2023), image segmentation (Xu et al. 2023a), and visual representation learning (Yu et al. 2024; Weng et al. 2024). Thus, we aim to utilize the strong capabilities of diffusion model to learn human-robot alignment and leverages the implicitly learned robot-dynamics features to guide the final policy learning.
@@ 4.1 Stage1 : Video Prediction Model (VPM)
We first explored the possibility of generating robot videos from human videos. Inspired by  (Hu 2024), our VPM explores a Spatial UNet (S-UNet in short) and a Spatial-Temporal UNet (ST-UNet in short), working collaboratively to learn from humans. In particular, the S-UNet extracts features from the robot arm, which is further fed into the ST-UNet for temporal modeling. VPM also contains a behavior extractors that estimate position and motion clues from human videos. Figure 3 gives an overview of our pipeline. Below, we introduce the architecture of VPM in detail.
@@
Spatial UNet. The S-UNet, consisting of 4 upsampling blocks and 4 downsampling blocks, extracts useful clues from the robotic arm. In particular, each block, known as the Spatial Layer (S-Layer in short), leverages self-attention to learn features tailored for robot arms, as well as cross-attention with pre-extracted CLIP embedding to incorporate semantic clues. The weights of the S-UNet are initialized from those of the Stable Diffusion (Rombach et al. 2022) to ease the learning process. The features derived by S-UNet provide a condition for predicting future frames.
@@
VAE Encoder and Decoder. Human2Robot builds upon Stable Diffusion (Rombach et al. 2022), which consists of an encoder and a decoder. The encoder turns images into latent embeddings for fast denoising, and the decoder maps the latents back to images. During training, both the encoder and decoder are kept frozen.
Training strategy. To generate higher-quality videos, the VPM training process consists of two stages. In the first stage, we focus on image generation to acquire basic generative capabilities. In the second stage, we train it to generate videos.
In the first stage, by taking the first frame of robot $o^{r}_{0}$, the first frame of human $o^{h}_{0}$ and future frame of human $o^{h}_{i}\ (i\in[0,T])$ as inputs to predict the future frame of robot $o^{r}_{i}$, where $T$ refer to the maximum length of video generation. In this stage, we train the S-UNet, the Behavior Extractors, and the ST-UNet without the temporal layers. The S-Layer in both the ST-UNet and S-UNet are instantiated using the powerful open-source Stable Diffusion model (SD)  (Rombach et al. 2022), while the Behavior Extractors are initialized with Gaussian weights, except for the final projection layer, which utilizes zero initialized convolution. In the second stage, we concentrate the training effort on the temporal layer to generate video. In this stage, we take a 30-frame segment of human video $o^{h}_{0:T}$ and the first frame of robot $o^{r}_{0}$ as inputs to predict the future robot observation $o^{r}_{0:T}$. The entire human video segment is first passed into Behavior Extractor, which is subsequently added to noise latent and fed to our ST-UNet. The weights of the VAE Encoder and Decoder, as well as the CLIP image encoder are frozen all the time.
@@
After the first stage of training, our VPM model is able to visually predict the corresponding robotic arm actions based on human movements. However, fully denoising an entire video is time-consuming, with most of the time spent on reconstructing image-level details, which are unrelated to manipulation. Increasingly, recent research (Hu et al. 2024; Wen et al. 2024; Zhu et al. 2025) has shown that the features of a generative model after a single denoising step already contain sufficient motion features to guide the action head’s planning. Inspired by these works, we treat the pretrained VPM as a vision encoder to extract the predicted robotic arm action information. This means adding intensity-invariant noise (t = K) to the human video and using the first denoised result as the feature. As shown in the Figure 5, we can see that the one-step denoising already contains a lot of action and position information.
Additionally, previous work (Weng et al. 2024; Xiang et al. 2023) has emphasized that the upsampling layers of diffusion models often contain more effective information. Therefore, we use the outputs of the VPM’s upsampling layers as features; more precisely, we use those from the first upsampling layer.
@@
Action Decoder. Inspired by recent video conditioned works (Hu et al. 2024; Luo and Lu 2025), our action decoder consists of two parts: the Video Former (Blattmann et al. 2023b) and the Diffusion Policy (Chi et al. 2023).
@@
To avoid the need to explicitly provide human demonstration videos for seen task, we use a k-nearest neighbors (KNN) approach to identify the most probable task for the current scene. We retrieve the human demonstration video corresponding to the closest matching features as the conditioning input to guide the task execution. Specifically, we use DINOv2 (Oquab et al. 2023) and CLIP (Radford et al. 2021) as feature extractors to capture features from the first frame of each robotic arm video in the training set. During prediction, we select the $n$ closest features based on the current environment, and the episode with the most frequent and closest match is chosen as the conditioning input, which is depicted in the Inference section of Figure 3.
@@ 5.1 Experimental Setups
  • Diffusion Policy (Chi et al. 2023): A action learning policy using action diffusers with CLIP language conditions. For simplicity, we refer to it as DP.
  • XSkill (Xu et al. 2023b): A human-video-conditioned policy through self-supervised learning.
  • Video Prediction Policy (Hu et al. 2024): A language-conditioned policy that uses video prediction for pretraining. For simplicity, we refer to it as VPP.
  • Action Decoder w. Human: Condition the action decoder of Human2Robot on human videos instead of the features extracted by VPM. The human videos are processed by ResNet18 (He et al. 2016) and used as input to action decoder.
@@ 5.3 Ablation Study
However, in our experiments, this approach produced highly jittery executions, and the model was insensitive to grasping motions in the human videos, frequently failing to complete grasps. As a result, the average success rate was only 23%. We attribute this to the complexity of human hand motions: without any prior for learning the correspondence, it is difficult to infer the correct mapping from human videos. By contrast, Human2Robot leverages video generation to learn the action correspondence, providing the downstream action decoder with a more reliable motion prior.
@@ 5.4 Generalization
Appearance Generalization. We tested the generalization ability on objects with different colors, materials, and textures. Human2Robot and VPP maintained their core capabilities without being affected, achieving success rates of 100% and 80%, respectively. However, XSkill could barely complete the tasks, performing only meaningless random movements, which demonstrates its poor generalization ability.
@@
Brand-New Task. Teaching a robot to write is a challenging problem for current learning-based approaches. The core issue lies in the vast number of characters, with writing each character being a brand-new task. We cannot teach the model to write each character individually, so it must possess strong generalization capabilities. To demonstrate the performance of Human2Robot in writing, we created a special dataset where each trajectory involves random movements on a desktop without writing any specific characters. We used this data to train the models to learn the correspondence between human and robots, and then during inference, we instructed the models to write brand-new characters, such as ”H” or ”R.” Experimental results show that only Human2Robot can learn the correspondence from meaningless data and successfully complete the task of writing specified character.
@@ Appendix A System Setup Details
Toward this goal, we build our H&R system to develop software and establish data collection protocols. We build the control system with OpenTeach (Iyer et al. 2024), connecting different hardware components to enable the operator to perform manual tasks and simultaneously monitor the robotic arm’s actions through the VR headset display. Specifically, we set up two similar environments for human operation and robotic arm operation respectively, as shown in Figure 7, and one camera is placed in each operating space. When the user wears the VR headset and performs hand movements, the headset transmits the coordinates of these movements to the backend, where they are retargeted to control the robotic arm, enabling synchronized actions. Section B illustrates more details about the retargeting process. Here, one of the biggest challenges lies in the inconsistency of the scale and the orientation between the two coordinate systems. In the following, we mainly illustrate our implementation for achieving the coordinate alignment.
@@ B.1 Coordinate System Alignment
To synchronize the rotation of the robotic arm with the human hand, we establish a 3D coordinate system similar to the approach taken by OpenTeach(Iyer et al. 2024). This coordinate system enables us to compute the rotation matrix $M_{h}^{t}$ for the human hand at time $t$. Additionally, we derive the rotation matrix $M_{r}^{t}$ for the robotic arm at the same time $t$, using the robot’s state parameters such as pitch, yaw, and roll. The rotation matrix $M_{h}^{t}$ and $M_{r}^{t}$ at the time $t$ can be expressed in terms of the rotation matrices $M_{h}^{0}$ and $M_{r}^{0}$ at time $t=0$ as:

### v4

@@ Abstract
Distilling knowledge from human demonstrations is a promising way for robots to learn and act. Existing methods, which often rely on coarsely-aligned video pairs, are typically constrained to learning global or task-level features. As a result, they tend to neglect the fine-grained frame-level dynamics required for complex manipulation and generalization to novel tasks. We posit that this limitation stems from a vicious circle of inadequate datasets and the methods they inspire. To break this cycle, we propose a paradigm shift that treats fine-grained human-robot alignment as a conditional video generation problem. To this end, we first introduce H&R, a novel third-person dataset containing 2,600 episodes of precisely synchronized human and robot motions, collected using a VR teleoperation system. We then present Human2Robot, a framework designed to leverage this data. Human2Robot employs a Video Prediction Model to learn a rich and implicit representation of robot dynamics by generating robot videos from human input, which in turn guides a decoupled action decoder. Our real-world experiments demonstrate that this approach not only achieves high performance on seen tasks but also exhibits one-shot generalization to novel positions, objects, instances, and even new task categories.
@@ 1 Introduction
The ability to observe others, whether humans or animals, and acquire skills to solve new tasks is a fundamental reason why humans excel at tackling a wide range of complex challenges. To enable robots to assist with diverse real-world problems, it is essential that they develop a similar capability, particularly the ability to learn directly from human demonstrations. This has motivated extensive research (bahl2023affordances; srirama2024hrp; nair2022rm; wang2023mimicplay; bahl2022human; smith2019avid) on how to learn from human demonstrations effectively.
Despite recent progress, current approaches face a fundamental generalization gap. While they perform well on tasks encountered during training, they often fail entirely when presented with human demonstrations of unseen tasks. The prevailing paradigm relies on applying self-supervised (xu2023xskill) or contrastive learning (jain2024vid2robot) methods to coarsely aligned human-robot video pairs. However, the lack of fine-grained supervision in existing datasets fundamentally constrains what models can learn. Rather than uncovering detailed action-level correspondences, models are often limited to capturing global features or high-level task summaries. Consequently, many approaches extract holistic video representations using models such as the Perceiver Resampler (alayrac2022flamingo), which compresses entire clips into fixed-length embeddings. This severely limits their ability to model the nuanced frame-level temporal dynamics that are essential for generalization.
@@
In this paper, we propose that the key to unlocking fine-grained human-robot alignment lies in video generation (vdm; gen1; align; tuneavideo; text2videozero; simda; tu2024motioneditor). Instead of merely learning whether two videos are broadly similar or mapping task goals from the human domain to the robot domain, we aim to generate a corresponding robot video directly from a human demonstration. We posit that by training a model to predict the precise frame-by-frame evolution of a robot’s movements, it learns a far richer and more temporally coherent alignment. This generative process forces the model to internalize the intricate dynamics of manipulation, enabling it to understand how a task is done, not just what is done.
@@ 2 Related Work
Teleoperation. Recently, VR-based methodologies (iyer2024open; ding2024bunny) have attracted considerable attention due to their cost-effectiveness, efficiency and versatility. However, these methods focus on controlling the robot, whereas our goal of using VR is to capture perfectly aligned videos of humans and robots, which are essential for robotic imitation.
Learning from Human Videos. Researchers have recently been attempting to leverage existing human-centric video datasets to enhance robot policy learning (liu2018imitation; smith2020avid; chen2021learning; videodex; zeng2024learning; zhang2025vlabench). Researchers propose to learn representations from human videos to assist in task execution (xiao2022masked; wang2023mimicplay; majumdar2024searchartificialvisualcortex). However, these approaches need strong prior knowledge and struggle to transfer to robots (nair2022rm; bahl2023affordances; bahl2022human). Meanwhile, some human-video-conditioned methods  (gen2act; jain2024vid2robot; xu2023xskill) have focused on aligning representations across human and robot videos, they are neither efficient nor capable of generalization. In contrast, our approach uses paired data and diffusion models to achieve strong generalization capabilities.
Diffusion Models for Video Generation. Current research has achieved remarkable performance of video generation (vdm; gen1; align; tuneavideo; text2videozero; simda; xing2023svformer; xing2025aid; xing2023vidiff; tu2024motioneditor; tu2024motionfollower; tu2024stableanimator). Moreover, diffusion models for video stylization (ye2025stylemaster; liu2023stylecrafter) and human image animation (hu2024animate; tu2024motioneditor) have also shown its promising potential for visual and motion transfer. Moreover, diffusion model not only performs impressively in generative domains but also shows promising applications in visual understanding task. Some notable approaches involve using diffusion models for object detection (diffusiondet; zhang2025diffusionad), image segmentation (xu2023open), and visual representation learning (yu2024representation; weng2024genrec). Thus, we aim to utilize the strong capabilities of diffusion model to learn human-robot alignment and leverages the implicitly learned robot-dynamics features to guide the final policy learning.
@@ 4.1 Stage1 : Video Prediction Model (VPM)
We first explored the possibility of generating robot videos from human videos. Inspired by  (hu2024animate), our VPM explores a Spatial UNet (S-UNet in short) and a Spatial-Temporal UNet (ST-UNet in short), working collaboratively to learn from humans. In particular, the S-UNet extracts features from the robot arm, which is further fed into the ST-UNet for temporal modeling. VPM also contains a behavior extractors that estimate position and motion clues from human videos. Figure 3 gives an overview of our pipeline. Below, we introduce the architecture of VPM in detail.
@@
Spatial UNet. The S-UNet, consisting of 4 upsampling blocks and 4 downsampling blocks, extracts useful clues from the robotic arm. In particular, each block, known as the Spatial Layer (S-Layer in short), leverages self-attention to learn features tailored for robot arms, as well as cross-attention with pre-extracted CLIP embedding to incorporate semantic clues. The weights of the S-UNet are initialized from those of the Stable Diffusion (rombach2022high) to ease the learning process. The features derived by S-UNet provide a condition for predicting future frames.
@@
VAE Encoder and Decoder. Human2Robot builds upon Stable Diffusion (rombach2022high), which consists of an encoder and a decoder. The encoder turns images into latent embeddings for fast denoising, and the decoder maps the latents back to images. During training, both the encoder and decoder are kept frozen.
Training strategy. To generate higher-quality videos, the VPM training process consists of two stages. In the first stage, we focus on image generation to acquire basic generative capabilities. Next, we train it to generate videos.
In the first stage, by taking the first frame of robot $o^{r}_{0}$, the first frame of human $o^{h}_{0}$ and future frame of human $o^{h}_{i}\ (i\in[0,T])$ as inputs to predict the future frame of robot $o^{r}_{i}$, where $T$ refer to the maximum length of video generation. In this stage, we train the S-UNet, the Behavior Extractors, and the ST-UNet without the temporal layers. The S-Layer in both the ST-UNet and S-UNet are instantiated using the powerful open-source Stable Diffusion model (SD)  (rombach2022high), while the Behavior Extractors are initialized with Gaussian weights, except for the final projection layer, which utilizes zero initialized convolution. In the second stage, we concentrate the training effort on the temporal layer to generate video. In this stage, we take a 30-frame segment of human video $o^{h}_{0:T}$ and the first frame of robot $o^{r}_{0}$ as inputs to predict the future robot observation $o^{r}_{0:T}$. The entire human video segment is first passed into Behavior Extractor, which is subsequently added to noise latent and fed to our ST-UNet. The weights of the VAE Encoder and Decoder, as well as the CLIP image encoder are frozen all the time.
@@
After the first stage of training, our VPM model is able to visually predict the corresponding robotic arm actions based on human movements. However, fully denoising an entire video is time-consuming, with most of the time spent on reconstructing image-level details, which are unrelated to manipulation. Increasingly, recent research (hu2024video; wen2024vidman; zhu2025unified) has shown that the features of a generative model after a single denoising step already contain sufficient motion features to guide the action head’s planning. Inspired by these works, we treat the pretrained VPM as a vision encoder to extract the predicted robotic arm action information. This means adding intensity-invariant noise (t = K) to the human video and using the first denoised result as the feature. As shown in the Figure 5, we can see that the one-step denoising already contains a lot of action and position information.
Additionally, previous work (weng2024genrec; xiang2023denoising) has emphasized that the upsampling layers of diffusion models often contain more effective information. Therefore, we use the outputs of the VPM’s upsampling layers as features; more precisely, we use those from the first upsampling layer.
@@
Action Decoder. Inspired by recent video conditioned works (hu2024video; luo2025learning), our action decoder consists of two parts: the Video Former (blattmann2023align) and the Diffusion Policy (chi2023diffusion).
@@
To avoid the need to explicitly provide human demonstration videos for seen task, we use a k-nearest neighbors (KNN) approach to identify the most probable task for the current scene. We retrieve the human demonstration video corresponding to the closest matching features as the conditioning input to guide the task execution. Specifically, we use DINOv2 (oquab2023dinov2) and CLIP (radford2021learning) as feature extractors to capture features from the first frame of each robotic arm video in the training set. During prediction, we select the $n$ closest features based on the current environment, and the episode with the most frequent and closest match is chosen as the conditioning input, which is depicted in the Inference section of Figure 3.
@@ 5.1 Experimental Setups
  • Diffusion Policy (chi2023diffusion): A action learning policy using action diffusers with CLIP language conditions. For simplicity, we refer to it as DP.
  • XSkill (xu2023xskill): A human-video-conditioned policy through self-supervised learning.
  • Video Prediction Policy (hu2024video): A language-conditioned policy that uses video prediction for pretraining. For simplicity, we refer to it as VPP.
  • Action Decoder w. Human: Condition the action decoder of Human2Robot on human videos instead of the features extracted by VPM. The human videos are processed by ResNet18 (he2016deep) and used as input to action decoder.
@@ 5.3 Ablation Study
However, in our experiments, this approach produced highly jittery executions, and the model was insensitive to grasping motions in the human videos, frequently failing to complete grasps. As a result, the average success rate was only 23%. We attribute this to the complexity of human motions: without any prior for learning the correspondence, it is difficult to infer the correct mapping from human videos. By contrast, Human2Robot leverages video generation to learn the action correspondence, providing the downstream action decoder with a more reliable motion prior.
@@ 5.4 Generalization
Appearance Generalization. We tested the generalization ability on objects with different colors, materials, and textures. Human2Robot and VPP maintained their core capabilities without being affected, achieving success rates of 100% and 80%, respectively. However, XSkill could barely complete the tasks, showing limited generalization.
@@
Brand-New Task. Teaching a robot to write is a challenging problem for current learning-based approaches. The core issue lies in the vast number of characters, with writing each character being a brand-new task. We cannot teach the model to write each character individually, so it must possess strong generalization capabilities. To demonstrate the performance of Human2Robot in writing, we created a special dataset where each trajectory involves random movements on a desktop without writing any specific characters. We used this data to train the models to learn the correspondence between human and robots, and then during inference, we instructed the models to write brand-new characters, such as ”H” or ”R.” Experimental results show that only Human2Robot can learn the correspondence from meaningless data and complete the task of writing character.
@@ 6 Conclusion
Acknowledgments.
This work was supported by the Science and Technology Commission of Shanghai Municipality(No. 24511103100).
@@ Appendix A System Setup Details
Toward this goal, we build our H&R system to develop software and establish data collection protocols. We build the control system with OpenTeach (iyer2024open), connecting different hardware components to enable the operator to perform manual tasks and simultaneously monitor the robotic arm’s actions through the VR headset display. Specifically, we set up two similar environments for human operation and robotic arm operation respectively, as shown in Figure 7, and one camera is placed in each operating space. When the user wears the VR headset and performs hand movements, the headset transmits the coordinates of these movements to the backend, where they are retargeted to control the robotic arm, enabling synchronized actions. Section B illustrates more details about the retargeting process. Here, one of the biggest challenges lies in the inconsistency of the scale and the orientation between the two coordinate systems. In the following, we mainly illustrate our implementation for achieving the coordinate alignment.
@@ B.1 Coordinate System Alignment
To synchronize the rotation of the robotic arm with the human hand, we establish a 3D coordinate system similar to the approach taken by OpenTeach(iyer2024open). This coordinate system enables us to compute the rotation matrix $M_{h}^{t}$ for the human hand at time $t$. Additionally, we derive the rotation matrix $M_{r}^{t}$ for the robotic arm at the same time $t$, using the robot’s state parameters such as pitch, yaw, and roll. The rotation matrix $M_{h}^{t}$ and $M_{r}^{t}$ at the time $t$ can be expressed in terms of the rotation matrices $M_{h}^{0}$ and $M_{r}^{0}$ at time $t=0$ as: